The Down and Out Co. just issued a dividend of $2.76 per share on its common stock. The company Is expected to maintain a constant 5 percent growth rate In its dividends Indefinitely. If the stock sells for $50 a share, what is the companys cost of equity? (Do not round your intermediate calculations.) 11.34% О 10.26% o 5.9% o 10.8% О 10.52%
Answer #1

Answer is 10.80%

Recent Dividend, D0 = $2.76
Growth Rate, g = 5%
Current Price, P0 = $50.00

Expected Dividend, D1 = D0 * (1 + g)
Expected Dividend, D1 = $2.76 * 1.05
Expected Dividend, D1 = $2.8980

Cost of Equity = D1 / P0 + g
Cost of Equity = $2.898 / $50.00 + 0.05
Cost of Equity = 0.0580 + 0.0500
Cost of Equity = 0.1080 or 10.80%